# Hui Xiong2, Bin Yang1

### v1

@@ 4.2 Model Analysis
Ablation study To ascertain the impact of different modules within CATCH, we perform ablation studies focusing on the following components: (1) Substitute the channel correlation discovering mechanism with fixed Channel Strategies. (2) Delete one of the four optimization objectives separately. (3) Remove the patching operation during training process. (4) Replace the Scoring technique with others. (5) Replace the bi-level optimization process with a normal process to optimize the mask generator and model simultaneously. Table 3 illustrates the unique impact of each module. We have the following observations: 1) Compared to the Channel-Independent (CI) Strategy, considering the relationships between variables using Channel Dependent (CD) Strategy or random masking yields better results, with the random masking performing worse than the CD method. Ours outperforms the CD method, further demonstrating the effectiveness of the channel correlation discovering mechanism. 2) Removing any of the four optimization objectives leads to a decline in model performance, with the most significant drop occurring when the frequency loss is removed. This fully demonstrates the rationality and effectiveness of the four optimization objectives. 3) When the patching operation is removed during training and replaced with a window-based approach to model the relationships between variables, the model performance significantly decreases. This indicates that the patching operation captures fine-grained frequency information, which is more conducive to anomaly detection. 4) When replacing the combination of point-granularity temporal anomaly scores and patch-wise point-aligned frequency anomaly scores with the combination of point-granularity temporal anomaly scores and window-granularity frequency anomaly scores, or when using only one of them, the model performance decreases in both cases. This indicates our Scoring technique shows stronger sensitivity in detecting anomalies. 5) When using a normal optimization process, the model performance also decreases consistently, which provides empircal evdience for the bi-level optimization.
@@
Variations CICIDS CalIt2 GECCO MSL SMD Average
@@
Anomaly criterion visualization We show how CATCH works by visualizing different anomalies in Figure 5. This figure showcases CATCH’s performance across five anomaly categories (Lai et al., 2021) on the TODS dataset, with temporal and frequency scores displayed in the second row, and the final anomaly scores in the third row. For point anomalies (first and second columns), the temporal scores exhibit sharp increases at the true anomaly locations, dominating the total scores. In contrast, for subsequence anomalies (third, fourth, and fifth columns), frequency scores remain elevated across the entire anomaly interval, show strong sensitivity to the actual boundaries of subsequence anomalies and compensate for the insensitivity of the temporal scores, which provides empircal evidence for the Frequency-Enhanced Point-Granularity Scoring technique in Section 3.4. Consequently, each domain contributes uniquely, allowing the final anomaly scores to accurately capture both point and subsequence anomalies.
@@ 5 Conclusion
6 Reproducibility
The study meets reproducibility requirements. Specifically, the datasets are available for download in a standardized format from here, and the code can be browsed at https://anonymous.4open.science/r/CATCH-E535. It may take some time to download the datasets, please wait patiently.
@@ A.1 Datasets
In order to comprehensively evaluate the performance of CATCH, we evaluate 9 real-world datasets and 12 synthetic datasets which cover 8 domains. The anomaly ratio vary from 0.17% to 11.07%, the range of feature dimensions varies from 3 to 72, and the sequence length varies from 5,040 to 1,416,825. This substantial diversity of the datasets enables comprehensive studies of MTSAD methods. Table 4 lists statistics of the 21 multivariate time series.
@@ A.3 Implementation Details
The “Drop Last” issue is reported by several researchers (Qiu et al., 2024; Xu et al., 2024). That is, in some previous works evaluating the model on test set with drop-last=True setting may cause additional errors related to test batch size. In our experiment, to ensure fair comparison in the future, we set the drop last to False for all baselines to avoid this issue.
All experiments are conducted using PyTorch (Paszke et al., 2019) in Python 3.8 and execute on an NVIDIA Tesla-A800 GPU. The training process is guided by the L1 loss, employing the ADAM optimizer. Initially, the batch size is set to 32, with the option to reduce it by half (to a minimum of 8) in case of an Out-Of-Memory (OOM) situation. We do not use the “Drop Last” operation during testing. To ensure reproducibility and facilitate experimentation, datasets and code are available at: https://anonymous.4open.science/r/CATCH-E535.
A.4 Implementation Details of Scoring
@@
Algorithm 2 Frequency-Enhanced Point-Granularity Scoring
@@
A.5 Command used for generating the synthetic datasets
@@ Appendix C Full Experimental Results
The full MTSAD results are provided in the following section due to the space limitation of the main text. Tables 5,  6,  7,  8, and 9, show the (AUC-ROC, R-AUC-ROC, VUS-ROC), (AUC-PR, R-AUC-PR, VUS-PR), (Accuracy, Precision, Recall, F1-score), (Range-Recall, Range-Precision, Range-F1-score), (Affiliated-Precision, Affiliated-Recall, Affiliated-F1-score) results, respectively.

### v3

@@ 4.2 Model Analysis
Ablation study To ascertain the impact of different modules within CATCH, we perform ablation studies focusing on the following components: (1) Substitute the channel correlation discovering mechanism with fixed Channel Strategies. (2) Delete one of the four optimization objectives separately. (3) Remove the patching operation during training process. (4) Replace the Scoring technique with others. (5) Replace the bi-level optimization process with a normal process to optimize the Mask Generator and model simultaneously.
@@
Variations CICIDS CalIt2 GECCO MSL SMD Avg
@@
Table 3 illustrates the unique impact of each module. We have the following observations: 1) Compared to the Channel-Independent (CI) Strategy, considering the relationships between variables using Channel Dependent (CD) Strategy or random masking yields better results, with the random masking performing worse than the CD method. Ours outperforms the CD method, further demonstrating the effectiveness of the channel correlation discovering mechanism. 2) Removing any of the four optimization objectives leads to a decline in model performance, with the most significant drop occurring when the frequency loss is removed. This fully demonstrates the rationality and effectiveness of the four optimization objectives. 3) When the patching operation is removed during training and replaced with a window-based approach to model the relationships between variables, the model performance significantly decreases. This indicates that the patching operation captures fine-grained frequency information, which is more conducive to anomaly detection. 4) When replacing the combination of point-granularity temporal anomaly scores and patch-wise point-aligned frequency anomaly scores with the combination of point-granularity temporal anomaly scores and window-granularity frequency anomaly scores, or when using only one of them, the model performance decreases in both cases. This indicates our Scoring technique shows stronger sensitivity in detecting anomalies. 5) When using a normal optimization process, the model performance also decreases consistently, which provides empircal evdience for the bi-level optimization.
@@
Anomaly criterion visualization We show how CATCH works by visualizing different anomalies in Figure 5. This figure showcases CATCH’s performance across five anomaly categories (Lai et al., 2021) on the TODS dataset, with temporal and frequency scores displayed in the second row, and the final anomaly scores in the third row. For point anomalies (first and second columns), the temporal scores exhibit sharp increases at the true anomaly locations, dominating the total scores. In contrast, for subsequence anomalies (third, fourth, and fifth columns), frequency scores remain elevated across the entire anomaly interval, show strong sensitivity to the actual boundaries of subsequence anomalies and compensate for the insensitivity of the temporal scores. Consequently, each domain contributes uniquely, allowing the final anomaly scores to accurately capture both point and subsequence anomalies.
@@ A.1 Datasets
ASD Application Server 19 1.55 12,848 4,320 Application Server Dataset collected from a large Internet company
@@
In order to comprehensively evaluate the performance of CATCH, we evaluate 10 real-world datasets and 12 synthetic datasets which cover 9 domains. The anomaly ratio vary from 0.17% to 11.07%, the range of feature dimensions varies from 3 to 72, and the sequence length varies from 5,040 to 1,416,825. This substantial diversity of the datasets enables comprehensive studies of MTSAD methods. Table 4 lists statistics of the 22 multivariate time series.
@@ A.3 Implementation Details
The “Drop Last” issue is reported by several researchers (Qiu et al., 2024; Li et al., 2024; Qiu et al., 2025b). That is, in some previous works evaluating the model on test set with drop-last=True setting may cause additional errors related to test batch size. In our experiment, to ensure fair comparison in the future, we set the drop last to False for all baselines to avoid this issue.
All experiments are conducted using PyTorch (Paszke et al., 2019) in Python 3.8 and execute on an NVIDIA Tesla-A800 GPU. We employ the ADAM optimizer during training. Initially, the batch size is set to 32, with the option to reduce it by half (to a minimum of 8) in case of an Out-Of-Memory (OOM) situation. We do not use the “Drop Last” operation during testing. To ensure reproducibility and facilitate experimentation, datasets and code are available at: https://github.com/decisionintelligence/CATCH.
A.4 Model Hyperparameter Settings
For each baseline method, we strictly follow the hyperparameter configurations recommended in their original papers. Additionally, we conduct hyperparameter searches on multiple sets and select the optimal configurations based on these evaluations to ensure a comprehensive and fair assessment of each method’s performance.
The hyperparameters for the baseline methods are set as follows:
  • AutoEncoder: The architecture of Autoencoder is selected from (10, 3), (25, 10, 5), (50, 20, 10).
  • One-Class SVM: The RBF kernel is used. The upper bound on the fraction of training errors is set to be 0.5.
  • Isolaion Forest: The number of estimators is selected from 3, 5, 25, 50, 60, 70, 80, 90, 100, 110.
  • Principal Component Analysis: The number of principal components is the smaller of the number of samples and the number of features.
  • HBOS: The number of bins is selected from 3, 5, 10, 20, 30, 40, 50
  • ModernTCN: The dimension of hidden states is 256. The FFN ratio is 2. Patch size is set as 16 and patch stride is set as 8.
  • Anomaly Transformer, DualTF: The number of layers is selected form 1, 2, 3, the channel number of hidden states dmodel is 512, and the number of heads is 8. The loss function hyperparameter $lambda$ for balancing two parts is set as 3.
  • iTransformer, TimesNet, DCdetector, PatchTST: The number of blocks is selected form 1, 2, 3. The dmodel is selected form 8, 64, 128, 256, 512 and the number of heads is 8.
  • DLinear, NLinear: The length of the reconstructed sequence is set to 100.
  • TFAD: The number of layers in the TCN is selected from 5, 6, 7, 8, and the number of kernels is selected from 5, 7.
A.5 Implementation Details of Scoring
@@
Algorithm 2 Calculation of freq-score
@@
A.6 Command used for generating the synthetic datasets
@@ Appendix C Full Experimental Results
The full MTSAD results are provided in the following section due to the space limitation of the main text. Tables 56,  7,  8,  9, and 10, show the (Accuracy), (AUC-ROC, R-AUC-ROC, VUS-ROC), (AUC-PR, R-AUC-PR, VUS-PR), (Precision, Recall, F1-score), (Range-Recall, Range-Precision, Range-F1-score), (Affiliated-Precision, Affiliated-Recall, Affiliated-F1-score) results, respectively.
Table 5: Average Acc (Accuracy) measures for all datasets. The best results are highlighted in bold, and the second-best results are underlined.
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE TFAD Ocsvm IF PCA HBOS
CICIDS Acc 0.758 0.750 0.762 0.565 0.745 0.750 0.750 0.751 0.757 0.754 0.994 0.952 0.082 0.852 0.656 0.913
CalIt2 Acc 0.946 0.889 0.930 0.910 0.944 0.841 0.946 0.889 0.887 0.844 0.960 0.942 0.867 0.968 0.884 0.886
Credit Acc 0.981 0.981 0.951 0.688 0.847 0.716 0.981 0.980 0.981 0.980 0.994 0.992 0.893 0.996 0.868 0.847
GECCO Acc 0.984 0.984 0.981 0.612 0.983 0.979 0.984 0.989 0.981 0.987 0.888 0.985 0.055 0.989 0.240 0.590
Genesis Acc 0.992 0.965 0.986 0.970 0.931 0.991 0.991 0.987 0.987 0.972 0.987 0.991 0.424 0.983 0.888 0.771
MSL Acc 0.853 0.857 0.854 0.891 0.891 0.892 0.855 0.873 0.858 0.872 0.891 0.891 0.818 0.890 0.812 0.816
NYC Acc 0.978 0.972 0.797 0.976 0.977 0.977 0.977 0.976 0.976 0.977 0.974 0.977 0.029 0.974 0.904 0.895
PSM Acc 0.730 0.729 0.728 0.697 0.719 0.615 0.726 0.729 0.726 0.728 0.715 0.720 0.737 0.723 0.682 0.736
SMD Acc 0.918 0.931 0.914 0.863 0.952 0.774 0.931 0.934 0.943 0.943 0.959 0.954 0.865 0.958 0.860 0.841
ASD Acc 0.933 0.913 0.898 0.851 0.942 0.930 0.928 0.942 0.947 0.912 0.946 0.950 0.871 0.953 0.863 0.870
Contextual4.9 Acc 0.961 0.734 0.957 0.713 0.641 0.618 0.734 0.888 0.943 0.742 0.965 0.874 0.888 0.956 0.848 0.865
Contextual7.2 Acc 0.960 0.721 0.937 0.603 0.702 0.611 0.739 0.886 0.893 0.729 0.956 0.859 0.877 0.939 0.818 0.850
Global4.8 Acc 0.993 0.896 0.994 0.833 0.174 0.639 0.989 0.995 0.993 0.895 0.980 0.876 0.903 0.977 0.698 0.872
Global7.2 Acc 0.978 0.849 0.972 0.326 0.232 0.665 0.987 0.984 0.985 0.900 0.978 0.855 0.915 0.970 0.644 0.843
Seasonal4.8 Acc 0.988 0.861 0.973 0.956 0.956 0.952 0.974 0.974 0.974 0.964 0.959 0.947 0.890 0.957 0.855 0.863
Seasonal7.7 Acc 0.979 0.792 0.964 0.894 0.928 0.925 0.963 0.960 0.965 0.943 0.938 0.916 0.887 0.939 0.839 0.849
Shapelet4.9 Acc 0.969 0.820 0.966 0.942 0.950 0.946 0.967 0.953 0.968 0.947 0.950 0.947 0.879 0.952 0.851 0.855
Shapelet7.4 Acc 0.957 0.715 0.949 0.871 0.921 0.823 0.951 0.943 0.947 0.891 0.932 0.913 0.872 0.925 0.838 0.839
Mixture5.7 Acc 0.903 0.893 0.885 0.658 0.928 0.940 0.896 0.889 0.400 0.882 0.692 0.921 0.183 0.948 0.183 0.190
Mixture8.9 Acc 0.808 0.849 0.822 0.638 0.882 0.910 0.822 0.814 0.587 0.804 0.650 0.859 0.123 0.923 0.124 0.215
Trend4.8 Acc 0.838 0.873 0.891 0.811 0.941 0.946 0.913 0.891 0.590 0.812 0.655 0.947 0.064 0.955 0.065 0.222
Trend7.8 Acc 0.911 0.813 0.790 0.785 0.903 0.911 0.801 0.878 0.603 0.864 0.640 0.893 0.079 0.927 0.080 0.319
$1^{st}$ Count 7 0 0 0 0 1 1 2 0 0 3 1 1 7 0 0